OpenStudy (anonymous):

Factoring/grouping question? x^2+xy+5x+5y -------------- x+5 So far I have (x^2+xy)+(5x+5y) __________________ x+5 Based on a guide I found it says that now I have to find the GCF between the two groups, but what would it be? x? If so, then how do I do it?

OpenStudy (akashdeepdeb):

NUMERATOR: \[x^2 + xy + 5x + 5y\] \[x(x+y) + 5(x+y)\] [Now take it common] Like so.. \[(x+y)(x+5)\] DENOMINATOR: \[(x+5)\] So Finally!! :D \[(x+y)(x+5)/(x+5) = (x+y)\] Getting me? :D

OpenStudy (anonymous):

I feel like I'm almost getting it! I see how you got x(x+y)+5(x+y) but how did you get just (x+y)(x+5)? I dont understand that part? D:

OpenStudy (akashdeepdeb):

Good question! :) It is just like this! ab + cb? = b(a+c) So x(x+y) + 5(x+y)? = (x+5)(x+y) Now getting it? :)

OpenStudy (akashdeepdeb):

It is just like using the Distributive property but in an opposite way! :D

OpenStudy (anonymous):

ahh I'm sorry I still don't get it >.< I kinda see what you mean though. What I have written down on my paper is x(x+y)+5(x+y) meeeh

OpenStudy (akashdeepdeb):

Great! :D Now look it is just like Alright Do this question! Break a(b+c) using distributive property! :D

OpenStudy (akashdeepdeb):

No problem! :D

OpenStudy (anonymous):

It would be ab+ac

OpenStudy (akashdeepdeb):

Good now! If I tell you to factorize this how will you do that?

OpenStudy (anonymous):

a(b+c)?

OpenStudy (jdoe0001):

\(\bf x^2+xy+5x+5y \implies x(x+y) + 5(x+y)\\ x\color{red}{(x+y)} + 5\color{red}{(x+y)} \textit{ common factor to } (x+5)\color{red}{(x+y)}\)

OpenStudy (jdoe0001):

see the common factor?

OpenStudy (anonymous):

I do!

OpenStudy (akashdeepdeb):

Exactly now take that COMMON And write like a(b+c) Tke the (x+y) as an 'a' So x(x+y) + 5(x+y) = xa + 5a = a(x+5) or (x+y)(x+5)!! :D Get it now? :D

OpenStudy (anonymous):

Oh! It all clicked! The answer was right! thanks for your time :D

OpenStudy (anonymous):

If I fully simplified it, it would be just x+y right?

OpenStudy (akashdeepdeb):

:) Yes! :D